मराठी
सी.आई.एस.सी.ई.आयसीएसई ICSE Class 7

Suppose S is the Event that Will Happen Tomorrow and P(S) = 0.03. Find P(S’)

Advertisements
Advertisements

प्रश्न

Suppose S is the event that will happen tomorrow and P(S) = 0.03. Find P(S’)

बेरीज
Advertisements

उत्तर

Given, P(S) = 0.03
P(S’) = 1 – P(S)
P(S’) = 1 – 0.03 ............[∵ P(S) + P(S’) = 1]
P(S’) = 0.97

shaalaa.com
Concept for Feel of Probability Using Data Through Experiments.
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 25: Probability - Exercise 22 (B)

APPEARS IN

संबंधित प्रश्‍न

A dice is thrown 20 times and the outcomes are noted as shown below:

Outcomes 1 2 3 4 5 6
No. of times 2 3 4 4 3 4

Now a dice is thrown at random, find the probability of getting:


In a cricket match, a batsman hits a boundary 12 times out of 80 balls he plays, further, if he plays one ball more, what will be the probability that:
(i) he hits a boundary
(ii) he does not hit a boundary


Two coins are tossed simultaneously 100 times and the outcomes are as given below:

Outcomes Two heads
(H, H)
Exactly one head
(H T or T H)
No head
(T T)
No. of times 21 55 24

If the same pair of coins is tossed again at random, find the probability of getting:
(i) two heads
(ii) exactly one head
(iii) no head.


A bag contains 4 white and 6 black balls,- all of the same shape and same size. A ball is drawn from the bag without looking into the bag. Find the probability that the ball drawn is:
(i) a black ball
(ii) a white ball
(iii) not a black ball


Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 50


Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 80


Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 40


Five Students A, B, C, D, and E are competing in a long-distance race. Each student’s probability of winning the race is given below:
A → 20%, B → 22%, C → 7%, D → 15% and E → 36%
(i) Who is most likely to win the race?
(ii) Who is least likely to win the race?
(iii) Find the sum of probabilities given.
(iv) Find the probability that either A or D will win the race.
(v) Let S be the event that B will win the race.
(a) Find P(S)
(b) State, in words, the complementary event S’.
(c) Find P(S’)


A Ticket is randomly selected from a basket containing 3 green, 4 yellow, and 5 blue tickets. Determine the probability of getting:
(i) a green ticket
(ii) a green or yellow ticket.
(iii) an orange ticket.


A box contains 3 yellow, 4 green, and 8 blue tickets. A ticket is chosen at random. Find the probability that the ticket is:
(i) yellow
(ii) green
(iii) blue
(iv) red
(v) not yellow


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×